# Deep Generative Classification of Blood Cell Morphology

### v1

@@ 1.7 Explainability
Figure 6: Counterfactual heatmaps for cell type transitions in the PBC dataset. The diagonal shows original images of each cell type, serving as the source. Each off-diagonal element in the same column represents a counterfactual heatmap ($\boldsymbol{H}_{c}$) showing the transition from the diagonal element (source) to the cell type of that row (target). Darker areas in the heatmaps indicate features to be added, whilst lighter areas suggest features to be removed to achieve the target cell type.
2 Discussion
This study introduces not only a novel model for haematological cell classification but, more crucially, a fundamental modelling approach and a principled evaluative framework. Both our approach and evaluative framework are grounded in the recognition of the inherent complexity of the target system and the constraints of applicable investigational and data regimes in medical imaging.
The characteristics of haematological cell classification mandate generative modelling, either as a preliminary to discriminative modelling or combined, as in our generative classifier, CytoDiffusion. This necessity arises from the fundamental inadequacy of learning a decision boundary in isolation. Such an approach will always be insufficient where the underlying distributions are as complex as those encountered in haematological cell morphology. Generative modelling, by contrast, can capture the full richness of these complex distributions.
Our empirical results manifest the failures of discriminative models, even state-of-the-art architectures trained on large-scale datasets, when subjected to an ecologically valid evaluative framework. This underscores the critical importance of assessing performance within this broader, more realistic context.
CytoDiffusion demonstrates substantial improvements over existing methods, without relying on careful hyperparameter tuning or data-specific architectural refinements. Importantly, these results likely represent a lower bound on potential performance, rather than a ceiling, highlighting the promise of this approach for further refinement and optimisation.
A key strength of our methodology is its recognition that the objective in medical imaging tasks like haematological cell classification is to achieve superhuman performance. This is attainable only by eliciting structure in the data that transcends expert labelling. While we have not fully exploited the representation learning capabilities of generative models in this work, there is significant potential for further advances, possibly through the use of generative modelling architectures with compact latents. Notably, CytoDiffusion’s confidence measure already demonstrates superhuman metacognitive capabilities, offering a more nuanced and reliable assessment of uncertainty than human experts.
Generative models also enable the assurance of equity in medical diagnostics. Although not implemented in this study, conditioning on minority characteristics would allow for legibility of differences in the model’s perception across diverse demographic groups. This capability is crucial for both detecting potential biases and enabling targeted remedial actions, ensuring fair and equitable application of AI in healthcare.
@@ 2 Discussion
Our approach is not without limitations. The inference process of our CytoDiffusion is computationally expensive, scaling with the number of classes. However, this is less problematic in the medical domain, where datasets typically have far fewer classes than general image classification tasks like ImageNet. Moreover, CytoDiffusion’s adaptive allocation of computing resources, dedicating more effort to challenging images, makes it more effective.
@@
In conclusion, this work establishes a new paradigm for medical image analysis, particularly in haematology. By embracing the inherent complexities of biological systems and the limitations of current data regimes, our generative approach with CytoDiffusion and comprehensive evaluation framework pave the way for more robust, interpretable, and equitable AI systems in healthcare. Future work should focus on refining these methods, exploring their applicability to other medical imaging domains, and further enhancing their computational efficiency without compromising performance.
3 Methods
3.1 Latent Diffusion Models
@@
Generative models have seen significant advancements in recent years, with diffusion models emerging as a leading approach for creating high-quality data samples [47, 48, 49, 50, 51]. These models draw inspiration from thermodynamic diffusion processes [52] and have demonstrated impressive results in various tasks, such as image generation, inpainting, and super-resolution [53, 47, 54, 55, 56].
Diffusion models operate on the principle of defining a forward diffusion process, which gradually adds noise to the data, converting it into a noise-like distribution. The model then learns a reverse process to denoise the data, effectively reconstructing the original data distribution [52, 49].
In latent diffusion models, data $\boldsymbol{x}$ is first encoded into a latent space using an encoder, typically a Variational Autoencoder (VAE), resulting in latent variables $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})$ with $\boldsymbol{z}\sim q(\boldsymbol{z})$ [57, 47, 58]. The forward and reverse diffusion processes are subsequently applied to these latent variables.
3.1.1 Forward Diffusion Process
@@
3.1.2 Reverse Diffusion Process
@@
The training objective for diffusion models is based on the variational lower bound (VLB) of the negative log-likelihood of the data. However, it was shown in [49] that a simpler objective could be used, which measures the difference between the added noise and the noise predicted by the model:
$\mathcal{L}_{\text{diffusion}}=\mathbb{E}_{\boldsymbol{\epsilon},t}\left[w_{t}\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t},t,c)\right\|^{2}_{2}\right]\,,$ (5)
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{I})$, and $t\sim\mathcal{U}[1,T]$ [49]. In our experiments, we use this objective, taking $w_{t}=\min\{\frac{5}{\text{SNR}(t)},1\}$ [59].
3.2 Diffusion Classifiers
Recent studies have shown that diffusion models can also be utilised as classifiers [18, 19, 20, 17]. Given an input image $\boldsymbol{x}$, we want to predict the most probable class $\hat{c}$. This can be formalised as finding the class $\hat{c}$ that maximises the posterior probability $p(c=c_{k}|\boldsymbol{x})$. Using Bayes’ theorem, this is equivalent to
with the weight $w_{t}$ as described below. To achieve this, we randomly pick a time step $t$ and noise tensor $\boldsymbol{\epsilon}$ and calculate the error
for all class labels $c$. These error values are then normalised using an appropriate weighting function. This normalisation aligns the errors for different $t$ and $\boldsymbol{\epsilon}$ to a common scale. Our empirical tests showed that the most effective weighting function was the reciprocal of a fourth-degree polynomial fitted to the median of the standard deviation of the squared errors. Further details on the weighting functions tested are available in the supplementary material A. After normalisation, the results are stored. The process then repeats with newly sampled values of $t$ and $\boldsymbol{\epsilon}$. For an intuitive explanation of how our model makes its predictions, we refer to Figure 10 in supplementary material.
Using a procedure similar to [19], we repeatedly gather new sets of errors for each candidate class. Classes that are less likely to be the lowest-error (and thus predicted) class as more samples are taken are progressively eliminated. This is achieved using a paired Student’s $t$-test, which compares the errors across classes. However, given that the errors do not perfectly follow the standard assumptions of a Student’s $t$-test, we employ a small p-value of $2\times 10^{-3}$ to ensure robustness. This iterative procedure continues until there is only one class left or a maximum number of iterations is reached. Unless stated otherwise, we use a minimum of 20 iterations and a maximum of 2000 iterations in our experiments.
3.3 General Training Setup
Unless otherwise specified, we employed the following training configuration for all experiments. We used Stable Diffusion 1.5 [47] as our base model. For class conditioning, we bypassed the tokeniser and text encoder, directly feeding the model with one-hot encoded vectors for each class, replicated vertically and padded horizontally to match the expected matrix of $77\times 768$ dimensions. We utilised a batch size of 10, a learning rate of $10^{-5}$ with linear warmup over 1000 steps, and trained on an A100-80GB GPU.
We applied various data augmentations, including random diagonal flips, 360-degree rotation, colour jitter (brightness=0.25, contrast=0.25, saturation=0.25, hue=0.125), and mixup [60] ($\alpha=0.3$) applied to the conditioning instead of the target. Additionally, we employed torchvision.transforms.RandAugment with default parameters. We used an AdamW optimiser ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, weight decay 0.01), mixed precision training (fp16), and an exponential moving average of 0.9999. All images were resized to $360\times 360$ pixels.
During inference, we applied the same data augmentations used in training, excluding torchvision.transforms.RandAugment and mixup. Leveraging the fact that white blood cells are typically centred in the image, and to mitigate the effects of augmentations that corrupt the outer parts of the image, we calculated the inference error only for pixels within a radius of 20 from the centre of the image in our latent space, which has dimensions of $45\times 45\times 4$.

### v2

@@ Results
Figure 6: Counterfactual heatmaps for cell type transitions in the PBC dataset. The diagonal shows original images of each cell type, serving as the source. Each off-diagonal element in the same column represents a counterfactual heatmap ($\boldsymbol{H}_{c}$) showing the transition from the diagonal element (source) to the cell type of that row (target). Areas in the heatmap with colours that deviate significantly from the background indicate regions where there is large error in the latent space between the two classes.
Discussion
This study introduces not only the novel CytoDiffusion model for haematological cell image classification but also a principled evaluative framework. Our approach is motivated by the desire to achieve a model with superhuman fidelity, flexibility and metacognitive awareness which can capture the distribution of all possible morphological appearances. These ambitions are attainable only by eliciting structure in the data beyond that which can be obtained simply from expert labelling of images. Both our methodology and evaluative framework are grounded in the recognition of the inherent complexity of the target system along with the inherent constraints for AI-modelling with medical imaging data, e.g. relatively small dataset regimes and instrumental variation reflected in images. In addition, we argue that medical imaging models demand a holistic evaluation framework that includes not only simple performance evaluation but also the model’s (a) robustness to domain shift, (b) ability to detect anomalies and (c) performance in low data availability scenarios, (d) ability to quantify its uncertainty accurately and (e) ability to provide interpretable outputs describing the decision making process.
Robustness of a classification model to domain shift, i.e. its ability to generalise across different imaging conditions, is crucial for its practical application in clinical settings. This generalisation capability is particularly important in haematology, where variations in microscope types, camera systems, and staining techniques are common across different laboratories and hospitals. A model that performs well only under specific imaging conditions would have limited utility in real-world clinical practice, potentially leading to inconsistent or unreliable diagnoses when deployed in new environments. Additionally, the ability to identify rare or unexpected cell types is crucial in clinical scenarios, where the detection of abnormal cells can have significant diagnostic implications. We propose that future models report performance scores for similar anomaly detection tasks, providing a standardised approach for assessing a model’s ability to flag potentially abnormal findings in medical imaging. Finally, the ability of a model to perform well with limited training data is crucial in medical imaging contexts, where large, well-annotated datasets are often challenging to obtain. We propose that new models in this domain should be evaluated on their performance in low-data regimes to assess their efficiency. Our results demonstrate the failures of discriminative models, even state-of-the-art architectures trained on large-scale datasets, when subjected to this more holistic evaluative framework whilst CytoDiffusion demonstrates substantial improvements, without relying on careful hyperparameter tuning or data-specific architectural refinements.
Additionally, in any classification task, whether performed by a human or an ML algorithm, it is highly informative to understand the uncertainty in the final decision, a surrogate for how difficult the sample is to classify. This is particularly true for clinical data where the classification result can inform an intervention or treatment decision. Currently, a method for optimal evaluation of model uncertainty is not established in the ML literature. We introduce a novel framework for evaluating an agent’s uncertainty–human or machine–based on the expected structure of the purely aleatoric uncertainty of an ideal psychophysical observer. This approach allows us to quantify model uncertainty as departure from the ideal psychometric function relating purely aleatoric uncertainty to fidelity. We demonstrate that CytoDiffusion produces superior estimates of uncertainty than human experts themselves.
Moreover, a key requirement of deploying machine learning models in clinical scenarios is their ability to provide interpretable results. CytoDiffusion addresses this need through the generation of counterfactual heatmaps, which highlight the regions of an image that would need to change for it to be classified as a different cell type, giving valuable insights into CytoDiffusion’s decision-making process, offering a detailed visualisation of the morphological distinctions the model identifies between cell types. This granular view of the model’s reasoning enhances our understanding of its classificatory mechanisms, potentially contributing to the development of more transparent and interpretable models across various domains of medical image analysis. A distinctive feature of our approach is that these heatmaps are the direct output of our model, requiring no post-processing or additional colour mapping. This direct output capability preserves the integrity of the model’s decision-making rationale, offering an unfiltered view of its reasoning process.
@@ Discussion
Our approach is not without limitations. The inference process of CytoDiffusion is computationally expensive, scaling with the number of classes. However, this is less problematic in the medical domain, where datasets typically have far fewer classes than general image classification tasks like ImageNet. Moreover, CytoDiffusion’s adaptive allocation of computing resources, dedicating more effort to challenging images, makes it more effective.
There are many possible extensions to this study. Firstly, we have not fully exploited the representation learning capabilities of generative models, where there is significant potential for further advances, possibly through e.g. the use of generative modelling architectures with compact latents. Additionally, generative models also enable the assurance of equity in medical diagnostics. Although not implemented in this study, conditioning on minority characteristics would allow for legibility of differences in the model’s perception across diverse demographic groups. This capability is crucial for both detecting potential biases and enabling targeted remedial actions, ensuring fair and equitable application of AI in healthcare.
@@
In conclusion, this work establishes a new paradigm for medical image analysis, particularly in haematology. By embracing the inherent complexities of biological systems, and the limitations of current data regimes, our generative approach with CytoDiffusion and comprehensive evaluation framework pave the way for more robust, interpretable, and equitable AI systems in healthcare. Future work should focus on refining these methods, exploring their applicability to other medical imaging domains, and further enhancing their computational efficiency without compromising performance.
Methods
Diffusion Classifiers
Recent studies have shown that diffusion models can also be utilised as classifiers [15, 16, 17, 14]. We use a latent diffusion model as described in Supplementary Material A. Given an input image $\boldsymbol{x}$, we want to predict the most probable class $\hat{c}$. This can be formalised as finding the class $\hat{c}$ that maximises the posterior probability $p(c=c_{k}|\boldsymbol{x})$. Using Bayes’ theorem, this is equivalent to
with the weight $w_{t}$ discussed in Supplementary Material A. To achieve this, we randomly pick a time step $t$ and noise $\boldsymbol{\epsilon}$ and calculate the error
for all class labels $c$. These error values are then normalised using $w_{t}$ and the results are stored. The process then repeats with newly sampled values of $t$ and $\boldsymbol{\epsilon}$. For an intuitive explanation of how our model makes its predictions, we refer to Supplementary Figure 10.
Using a procedure similar to [16], we repeatedly gather new sets of errors for each candidate class. Classes that are less likely to have the lowest-error are progressively eliminated. This is achieved using a paired Student’s $t$-test, which compares the errors across classes. However, given that the errors do not perfectly follow the standard assumptions of a Student’s $t$-test, we employ a small p-value of $2\times 10^{-3}$ to ensure robustness. This iterative procedure continues until there is only one class left or a maximum number of iterations is reached. Unless stated otherwise, we use a minimum of 20 iterations and a maximum of 2000 iterations in our experiments. For an analysis of different p-values tested, as well as various minimum and maximum numbers of iterations, we refer to Supplementary Figure 7.
General Training Setup
Unless otherwise specified, we employed the following training configuration for all experiments. We used Stable Diffusion 1.5 [38] as our base model. For class conditioning, we bypassed the tokeniser and text encoder, directly feeding the model with one-hot encoded vectors for each class, replicated vertically and padded horizontally to match the expected matrix of $77\times 768$ dimensions. We utilised a batch size of 10, a learning rate of $10^{-5}$ with linear warmup over 1000 steps, and trained on an A100-80GB GPU. Details of the training and inference parameters are in the Supplementary Materials B.
Datasets
We utilised multiple datasets, described in Table 3, including four which are publicly available and one custom dataset, CytoData, to develop and evaluate our diffusion classifier for haematological cell image classification. CytoData, available at [insert upon publication], is an anonymised dataset consisting of 559,808 single cell images from 2,904 blood smear slides obtained from Addenbrooke’s Hospital in Cambridge, UK, with a labelled subset of 5,000 images across 10 classes. These images were created using CellaVision, a specialised imaging technology for cellular analysis. The labelling strategy is described in the Supplementary Materials C. Notably, when labelling CytoData, we included an artefact class, addressing a critical challenge in clinical applications, as blood smear slides often contain artefacts which may be mistaken for cells by deep learning-based cell detection models. By explicitly modelling these artefacts, CytoData aims to enhance clinical applicability. A distinctive feature of CytoData is the inclusion of labeller confidence scores. This addition provides valuable information for various analyses beyond simple correlations.
Table 3: Summary of the datasets used in our experiments.
Dataset # Images # Classes Labeller Confidence Artefacts
[…]
LISC [40] 257 5
PBC [19] 17,092 8
Bodzas [35] 16,027 9
CytoData (ours) 5,000 $9+1$*
[…]
    Split into training (10,175 images), Test-A (4,339 images), and Test-B (2,119 images) sets.
    9 regular classes of blood cells plus 1 artefact class.
Authenticity Test
To assess the quality and authenticity of our fine-tuned diffusion model’s synthetic blood cell images, we conducted an authenticity test with expert haematologists. This evaluation was designed to determine whether the model could effectively capture the underlying distribution of blood cell images across various cell types, a capability that traditional discriminative models are not inherently required to possess. Additionally, we sought to assess the accuracy of the generated cell types. Further details are in Supplementary Materials D. Ten haematology specialists from our research group, with {34, 28, 25, 15, 10, 9, 6, 5, 5, 1} years of experience in blood microscopy, participated in the authenticity test. Participants were informed that half of the images presented to them would be real images from our dataset, while the other half would be synthetic images generated by our model. Each specialist was presented with the 288 images in a randomised order and asked to perform two tasks: (1) identify whether each image was synthetic or real and (2) classify each image into one of the nine designated blood cell types.
In-Domain Performance
To establish a baseline for our diffusion-based classifier, CytoDiffusion, we evaluated its performance on standard in-domain classification tasks using three datasets: our custom dataset, Raabin-WBC, and PBC. For the Raabin-WBC dataset, we utilised the predefined train split (90% for training, 10% for validation) and tested on the predefined Test-A dataset. For the PBC dataset, we employed an 80–10–10 split for train–validation–test. For our custom dataset, we employed a cross-validation approach where we divided the dataset into four folds and fine-tuned four separate models. For each fold used in training, we further split the data into 80% for training and 20% for validation and we trained each model for 60,000 iterations. For the Raabin-WBC and PBC datasets, we trained our model for 72,000 steps. To provide a basis for comparison, we also trained and evaluated EfficientNetV2-M and ViT-B/16 models. Given the significant class imbalances present in these datasets (Raabin-WBC Test-A ranges from 89 to 2,660 images per class, our custom dataset from 103 to 1,425, and PBC from 1,199 to 3,314), we report both overall accuracy and balanced accuracy. It is important to note that we have excluded some studies from our comparison due to methodological differences that could lead to unfair or misleading comparisons. Specifically, we are not comparing with papers that do not have a conventional train–validation–test split [41, 42], or those that do not test on the predefined test set [32, 43].
Uncertainty Measure
To evaluate the quality of CytoDiffusion’s uncertainty measure, we exploited the decomposability of uncertainty into model and aleatoric components. An ideal model, indeed any ideal agent, should contribute no uncertainty of its own, leaving aleatoric uncertainty as the sole residue. If so, then the uncertainty measure should reduce to the magnitude of the discriminative signal, and the relation between the uncertainty measure and model fidelity should conform to that of an ideal observer of a noisy signal. This allows us to exploit psychometric function modelling to quantify how close an agent’s uncertainty–machine or human–is to the aleatoric floor [44]. We utilised the cross-validated models trained on our custom dataset, as described in the “In-Domain Performance” section. For quantifying the model’s uncertainty, we calculated the difference between the two classes with the smallest error, rescaled in the interval $[0,1]$. To establish a measure of labeller uncertainty, we mapped the confidence levels provided by our expert haematologists to numerical values: 1.0 for High Confidence, 2/3 for Moderate Confidence, 1/3 for Low Confidence, and 0.0 for No Confidence. For images assessed by multiple labellers, we averaged their uncertainty scores.
Psychometric functions describe the relation between the performance of an observer and a (typically scalar) property of the observed [45]. The performance of interest is usually detection or classification, expressed as a function of signal strength on a monotonically increasing scale. Since an ideal model is as confident as the data allows, exhibiting purely aleatoric uncertainty, we can quantify the proximity of a model to that ideal by fitting a psychometric function with model confidence as the index of signal strength. A good measure of uncertainty should conform closely to an ideal observer, yielding a sigmoid curve rising from a chance guess rate, $\gamma$, where uncertainty is maximal, to a lapse rate $\lambda$, where uncertainty is minimal and any errors are not explicable by insufficient information. Further details are in the Supplementary Material E. For $\gamma$, we fix the parameter at 1/10, reflecting the 10 possible classes, and for $\lambda$ we use a beta-distribution with parameters (1,10). We report the threshold and width estimated parameters, and their posterior distributions, citing 95% credibility intervals.
Anomaly Detection
To evaluate our model’s capability for detecting anomalous cell types, we designed an experiment that simulates real-world scenarios where rare or previously unseen cell types might appear in clinical samples. This approach involved excluding specific abnormal cell classes during training and assessing the model’s ability to identify these classes during testing. We utilised two datasets for this experiment: the Bodzas dataset and the PBC dataset. For the Bodzas dataset, we merged the neutrophil segment and neutrophil band classes into a single neutrophil class and excluded the blasts class, which comprised both lymphoblasts and myeloblasts (5,036 images in total). From the PBC dataset, we excluded the erythroblast class (1,513 images). The data was split into 70% training, 10% validation, and 20% test sets, with all abnormal cells (blasts and erythroblasts) moved to their respective test sets. We employ an anomaly score inspired by [46, 47] to quantify the model’s confidence (details in Supplementary Material F). To visualise each model’s ability to distinguish between normal and abnormal cells, we generated KDE curves of the confidence differences for both groups. To quantify this ability, we calculated the AUC for each model and dataset.
Domain Shift
To evaluate our model’s robustness under domain shift conditions, we leveraged the Raabin-WBC dataset [39] and the LISC dataset [40], following methodologies outlined in previous studies [25, 24]. We utilised the predefined train split for training (90%) and validation (10%), and the Test-B split as one of our test sets. The Test-B split was created using a different microscope and camera type compared to the training and validation set, introducing a domain shift. Additionally, we employed the LISC dataset as a second test set, which was created using different microscope and camera types, as well as a different staining method, further increasing the domain shift challenge. To maintain consistency with previous studies [25, 24] and accommodate the lower resolution of LISC images, we resized all images to $224\times 224$ pixels. We used a batch size of 32 and trained for 22,000 steps on an NVIDIA RTX A5000 GPU. For comparison, we also fine-tuned and tested EfficientNetV2-M and ViT-B/16. To ensure reliable performance evaluation, we repeated the training and evaluation process five times for each of the models (our diffusion-based model, EfficientNetV2-M, and ViT-B/16).
Efficiency in Low-Data Regimes
To assess the efficiency of our model in learning from a small number of images per class, we designed an experiment using the PBC dataset [19]. We randomly selected 250 images per class for the training set, 836 images per class for the test set, and the remaining images for the validation set. From the training set, we randomly sampled subsets of 10, 20, 50, or 150 images per class to simulate low-data scenarios. We trained CytoDiffusion on an NVIDIA RTX A5000 GPU. For subsets of 10, 20, 50, and 150 images per class, we trained the model for 20,000, 40,000, 100,000, and 240,000 steps, and saved checkpoints every 1,000, 3,000, 7,000, and 20,000 steps, respectively. For each subset, we selected the checkpoint with the highest validation accuracy for testing. For comparison, we also trained and evaluated EfficientNetV2-M and ViT-B/16 models using the same data subsets. To account for variability in the subsampling and initialisations, we repeated each experiment five times. This involved resampling the specified number of images per class (10, 20, 50, and 150) and retraining all models for each repetition.
Explainability
One of the integral aspects of CytoDiffusion’s utility in clinical settings is its ability to provide explainable predictions. To achieve this, we employ a counterfactual heatmap approach, which elucidates what changes would be necessary for an image to be classified under a different specified class. This method is particularly beneficial for understanding model decisions in complex medical imaging tasks such as the classification of blood cell types. Initially, we calculate the difference between the original noise, $\boldsymbol{\epsilon}$, and the noise predicted by the model for each class condition $c$. This difference is recorded for all iterations, and the mean error for each condition $c$ is computed as $\Delta_{c}=\frac{1}{N}\sum_{n=1}^{N}\left(\boldsymbol{\epsilon}_{n}-\boldsymbol{\epsilon}_{\theta}(\boldsymbol{z}_{t_{n}},t_{n},c)\right)\,$ , where $N$ is the number of iterations. Subsequently, for each class condition $c$, we calculate the deviation from the condition with the minimum error, designated as $\Delta_{\hat{c}}$, where $\hat{c}$ is the predicted class. The adjusted $\delta_{c}$ is then $\delta_{c}=\Delta_{c}-\Delta_{\hat{c}}\,$. Finally, $\delta_{c}$ are decoded back to the pixel space using the VAE decoder to obtain the counterfactual heatmaps $\boldsymbol{H}_{c}=\mathcal{D}(\delta_{c})\,$, where $\mathcal{D}$ denotes the VAE decoder, and $\boldsymbol{H}_{c}$ represents the heatmap for condition $c$. These heatmaps visually represent modifications that would shift the image classification from the predicted class to the target class $c$, thereby providing a powerful tool for explaining and validating model predictions.
Competing interests. The authors declare no competing interests.
Data availability. CytoData is available at [insert upon publication]. All other datasets used in the analysis are publicly available as referenced.
Code availability. All code and trained models are available at https://github.com/CambridgeCIA/CytoDiffusion.
BloodCounts! Consortium. Martijn Schut, Folkert Asselbergs, Sujoy Kar, Suthesh Sivapalaratnam, Sophie Williams, Mickey Koh, Yvonne Henskens, Bart de Wit, Umberto D’Alessandro, Bubacarr Bah, Ousman Secka, Parashkev Nachev, Rajeev Gupta, Sara Trompeter, Nancy Boeckx, Christine van Laer, Gordon A. Awandare, Kwabena Sarpong, Lucas Amenga-Etego, Mathie Leers, Mirelle Huijskens, Samuel McDermott, Willem H. Ouwehand, James Rudd, Carola-Bibiane Schönlieb, Nicholas Gleadall and Michael Roberts.
Appendix A Latent Diffusion Models Setup
@@
Generative models have seen significant advancements in recent years, with diffusion models emerging as a leading approach for creating high-quality data samples [38, 48, 49, 50, 51]. These models draw inspiration from thermodynamic diffusion processes [52] and have demonstrated impressive results in various tasks, such as image generation, inpainting, and super-resolution [53, 38, 54, 55, 56]. Diffusion models operate on the principle of defining a forward diffusion process, which gradually adds noise to the data, converting it into a noise-like distribution. The model then learns a reverse process to denoise the data, effectively reconstructing the original data distribution [52, 49]. In latent diffusion models, data $\boldsymbol{x}$ is first encoded into a latent space using an encoder, typically a Variational Autoencoder (VAE), resulting in latent variables $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})$ with $\boldsymbol{z}\sim q(\boldsymbol{z})$ [57, 38, 58]. The forward and reverse diffusion processes are subsequently applied to these latent variables.
A.1 Forward Diffusion Process
@@
A.2 Reverse Diffusion Process
@@
The training objective for diffusion models is based on the variational lower bound (VLB) of the negative log-likelihood of the data. However, it was shown in [49] that a simpler objective could be used, which measures the difference between the true noise added and that predicted by the model
$\hat{c}=\operatorname*{argmin}_{c}\mathcal{L}_{\text{diffusion}}(c)\,,$ (8)
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{I})$, $t\sim\mathcal{U}[1,T]$ and $w_{t}=\min\{\frac{5}{\text{SNR}(t)},1\}$ is a time-dependent weighting function that balances the importance of different timesteps during training [59, 49].
Appendix B Training and Inference Parameters and Procedure
We applied various data augmentations, including random diagonal flips, random rotation (with angles uniformly sampled between 0 and 359 degrees), colour jitter (brightness = 0.25, contrast = 0.25, saturation = 0.25, hue = 0.125), and Mixup [60] ($\alpha=0.3$) applied to the conditioning instead of the target and RandAugment with default parameters [61]. We used an AdamW optimiser ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, weight decay 0.01), mixed precision training (fp16), and an exponential moving average of 0.9999. All images were resized to $360\times 360$ pixels.
During inference, we applied the same data augmentations used in training, excluding RandAugment and Mixup. Leveraging the fact that white blood cells are typically centred in our images, and to mitigate the effects of augmentations that corrupt the outer parts of the image, we calculated the inference error only for pixels within a radius of 20 from the centre of the image in our latent space, which has dimensions of $45\times 45\times 4$.